Calculate the strength and direction of the electric field  E  due to a point charge point charge of 5.0  nC   at a distance of 8.00 mm from the charge. 


1
Expert's answer
2021-04-28T09:56:06-0400

The electric field due to a point charge "q = 5n C = 5\times 10^{-9}C" at distance "r = 8mm = 0.008m" is given as follows:


"E = k\dfrac{q}{r^2}"

where "k = 9\times 10^{9}N\cdot m^2/C^2" is the Coulomb's constant. Thus, obtain:


"E = 9\times 10^{9}\dfrac{N\cdot m^2}{C^2}\cdot \dfrac{5\times 10^{-9}C}{(0.008m)^2} \approx 7\times 10^5N/C"

Answer. "7\times 10^5N/C".